OpenStudy (anonymous):

equation sets use digits:1 to 6 e.g.:6/3=5+2-4-1

OpenStudy (anonymous):

6(2)-3(1)=5+4

OpenStudy (anonymous):

any?

OpenStudy (anonymous):

\[(\sqrt{4})\div2=(5+3)-(6+1)\]

OpenStudy (anonymous):

(5-2)/3=6-4-1 5*3=4*2+6+1 6*3=(5+4)*2*1 6+5=4*2+3+1

OpenStudy (anonymous):

the last one should be 6+5=4*2 + 3*1

OpenStudy (anonymous):

* mean multiples right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much :) can you read my another question